# LoRACLR: Contrastive Adaptation for Customization of Diffusion Models

### v1

@@ 2 Related Work
Merging Multiple Concepts. Combining LoRAs for style and subject control remains an open research challenge. Current methods for multi-concept synthesis often face notable limitations. Weighted summation [32] is a simple approach but suffers from feature interference. Mix-of-Show [9] requires specialized Embedding-Decomposed LoRAs (ED-LoRAs) for each concept, which limits compatibility with standard LoRAs. ZipLoRA [35] can merge style and content LoRAs but struggles when multiple content LoRAs are needed. OMG [19] depends on off-the-shelf segmentation to isolate subjects, making its performance highly dependent on segmentation accuracy and the model’s ability to generate multiple objects.
Recent advancements have focused on merging multiple specialized models into a unified generative framework. Mix-of-Show [9] effectively merges models using ED-LoRAs but requires access to the original training data, which limits compatibility with community LoRAs, such as those available on platforms like civit.ai [3]. Orthogonal Adaptation [24] introduces constraints to separate attributes across LoRAs, reducing interference; however, it increases training complexity by directly modifying the fine-tuning process, which also requires access to the original data.
@@ 3.2 LoRACLR
The core of LoRACLR’s merging process is a contrastive loss objective designed to ensure compatibility across independently trained LoRA models within a unified model. Given a pre-trained LoRA $V_{i}$, we first create pairs of input and output features, denoted as $X_{i}$ and $Y_{i}$, respectively, see Fig. 2 (a). Meanwhile, predicted features $\hat{Y_{i}}$ represent the merged model’s output for the same input features, $X_{i}$, i.e., $\hat{Y_{i}}=(W+\Delta W)X_{i}$. The intuition behind our contrastive objective is as follows: positive pairs generated by the same LoRA models should attract, while negative pairs generated by the different LoRAs should repel, see Fig. 2 (b). The contrastive loss objective is defined as:
@@
where $d_{n,i}$ is the negative distance, computed as the minimum distance between the output features $Y_{i}$ for a given concept and the predicted features $\hat{Y}_{j}$ of unrelated concepts, with $j\neq i$. This contrastive objective keeps unrelated concepts distinct while aligning each concept’s output features with the merged model’s predicted features, enabling LoRACLR to perform coherent multi-concept synthesis with minimal interference.
Delta-Based Merging. LoRACLR uses an additive delta, $\Delta W$, to merge LoRA models without altering the base weights directly. Initialized to zero, we learn $\Delta W$ to adjust the pre-trained weights, preserving the integrity of each model while ensuring compatibility. An L2 regularization term is applied to $\Delta W$ to limit its magnitude, ensuring sparsity and minimal adjustments. The optimization objective combines the contrastive merging loss, $\mathcal{L}_{\text{contrastive}}$ and $\mathcal{L}_{\text{delta}}$:
@@ 4 Experiments
Figure 5: Quantitative Results on Number of Concepts. Text alignment, image alignment, and identity preservation scores as the number of merged concepts increases. Our method achieves high scores across all metrics, maintaining identity and prompt adherence. Dots represent the baseline metrics for each LoRA model before merging, serving as a reference for performance comparisons.
Implementation Details. In all our experiments, we use the Stable Diffusion model [29] with the ChilloutMix checkpoint [^1] https://huggingface.co/windwhinny/chilloutmix for its high-quality image generation capability. Our approach leverages pre-trained LoRA models, eliminating the need to train individual LoRA models from scratch and significantly reducing computational overhead. During the merging process, we use a learning rate of $1e^{-4}$, a margin parameter $m$ of 0.5 to control separation between positive and negative pairs during contrastive learning, and a regularization coefficient $\lambda_{\text{delta}}$ of 0.001 to ensure sparsity in the learned delta. All experiments, including model outputs, and processing, were conducted on an NVIDIA A100 GPU at Virginia Tech. Combining 12 concepts is highly efficient, typically taking about 5 minutes, which makes our approach scalable and practical for real-world applications that require seamless integration of multiple concepts. For our experiments, we utilize pre-trained single-concept LoRA and/or ED-LoRA models as the starting point.
Baselines. We evaluate our approach against several state-of-the-art baselines in multi subject customization, including: DB-LoRA [30], $\mathcal{P}+$ [39], Custom Diffusion [20], Mix-of-Show [9], and Orthogonal Adaptation [24]. Each method employs a different strategy for merging fine-tuned models. For DB-LoRA, Federated Averaging (FedAvg) is used to merge models. Custom Diffusion utilizes its optimization-based method for merging, while Mix-of-Show merges using gradient fusion. Orthogonal Adaptation introduces orthogonal transformations to mitigate concept interference during merging. In the case of $\mathcal{P}+$, no fine-tuning of model weights is performed; instead, merging is achieved by directly querying each concept’s token embedding.
Datasets & Metrics. For all evaluations, we use the same experimental setup and dataset proposed by [24], which includes 12 concept identities, each represented by 16 distinct images of the target concept across various contexts. Following prior work [24, 9], we evaluate our method using three key metrics: Text Alignment, Image Alignment, and Identity Alignment. Text Alignment measures text-image similarity using the CLIP [26] model to ensure that generated images match the input prompts. Image Alignment evaluates the similarity between generated and reference images in the CLIP feature space. Identity Alignment uses the ArcFace [4] model to assess how accurately the target human identity is preserved in the generated images.
4.1 Qualitative Results
We demonstrate the qualitative results of our method for both single and multiple subjects. Given a set of subjects, each represented by a personalized model—for instance, celebrities like Margot Robbie and Taylor Swift (refer to Fig. 3)—and a text prompt such as ‘<Margot> and <Taylor> in an ancient library with towering shelves,’ our objective is to generate a composite image that integrates these subjects according to the provided text prompt. Using 12 subjects identified by [24], we first merge them into a unified model through our novel contrastive-based objective, applied consistently across all experiments. Unlike methods that require individual fine-tuning for each concept [24], our approach can work with pre-trained models such as LoRA or ED-LoRA models.
Single Concepts. We first demonstrate our method’s ability to preserve individual identities. Figure 3 (bottom row) confirms that our approach maintains the integrity of each identity for single concepts. This capability extends to diverse settings. For example, the “<CONCEPT> Cyberpunk style’ shows subjects portrayed as game-like characters, as seen in the fourth images of the bottom row in Fig. 3.
Multiple Concepts. Figure 3 showcases images generated with our method using varying numbers of concepts—6, 5, 4, 3, and 2. These visuals demonstrate that our method not only accurately captures each individual identity but also generates composite images guided by the text prompts. Notably, our approach excels by utilizing a single merged model to generate a wide range of concepts across diverse prompts, eliminating the need to train separate models for different concept counts. We also note that the original LoRA models for individual subjects are trained on various images of celebrities who may have different hairstyles or colors. As a result, variations such as different hairstyles or colors can appear even within generations of the same celebrity model (e.g., different generations of a Taylor Swift LoRA might show varying hair colors or styles, reflecting her real-life changes). These variations are not inconsistencies in our model but rather reflect the diversity inherent in the original LoRA models. Importantly, despite these variations, our model maintains fidelity and identity to the celebrities’ faces and ensures that details like hairstyles do not mix across different subjects.
Qualitative Comparisons. In Fig. 4, we visually compare our method against several state-of-the-art approaches, including Orthogonal Adaptation, MoS, and $\mathcal{P}+$, across various concepts. Figure 4 illustrate that our method successfully preserves individual identities without crossovers, whereas other methods encounter issues. For instance, Orthogonal Adaptation [24] inadvertently transfers features, such as the hair of female characters, to others, while MoS [9] and $\mathcal{P}+$ [39] struggle with precise identity depiction. Additionally, while these models perform adequately with two or three concepts, their ability to accurately represent more concepts diminishes as the concept count increases. For instance, in scenes with six concepts, refer to Fig. 4 first section, methods other than LoRACLR fail to preserve the identity of Lionel Messi, highlighting their limitations as the number of concepts increases.
Style LoRA Integration To showcase the flexibility of our approach, we integrate style-specific LoRA models to generate scenes combining conceptual and stylistic elements, enabling outputs in styles like comic art or oil painting. As shown in Fig. 6, prompts such as “…in a city, in comic style” capture the vibrant aesthetic of comic art, while “…in a garden, holding flowers, in oil painting style” reflect the textured quality of oil painting. Complex scenes like “…in a castle, signing papers, in oil painting style” illustrate the model’s ability to maintain content coherence alongside stylistic adaptation. These results demonstrate our method’s efficiency in preserving content accuracy and achieving high stylistic fidelity, making it practical for creative and artistic workflows.
@@ 4.1 Qualitative Results
Figure 6: Style LoRA Integration. Our method combines styles like comic art and oil painting into multi-subject scenes, ensuring stylistic fidelity and content coherence, showcasing its flexibility.
Figure 7: Non-human subject generation. Our method effectively combines diverse concepts such as animals, objects (e.g., tables, chairs, vases), and monuments (e.g., pyramids, rocks) into cohesive and visually appealing scenes.
@@
Non-Human Examples. We present examples involving non-human concepts, including animals, objects, and monuments. Our model generalizes seamlessly to these scenarios while maintaining high-quality outputs and stylistic consistency. As shown in Fig. 7, our method effectively integrates diverse concepts into cohesive scenes. It excels with objects like tables, chairs, and vases, accurately capturing their distinct textures. Additionally, it demonstrates robustness with monumental elements such as pyramids and rocks, blending them seamlessly with other concepts. These results highlight the adaptability of our method in generating coherent and visually engaging compositions across a wide range of non-human concepts, including animals, objects, and landmarks. This versatility makes it suitable for both creative and practical applications, such as wildlife illustration, interior design, and architectural visualization.
@@ 4.2 Quantitative Results
Effect of # Concepts. Unlike other methods, which struggle to maintain text alignment, image alignment, and identity preservation as the number of combined concepts increases, LoRACLR maintains these metrics, see 5.
User Study. To supplement the findings in Tab. 1 and Fig. 5, we conduct a user study involving 50 participants on Prolific.com [25]. The study consisted of 40 questions where participants are shown reference images of individual concepts alongside composite images, with varying numbers of concepts, generated by our method and competing methods, presented in a randomized order. Participants are asked to evaluate each image pair based on: Identity Alignment: Given the reference images on the left, how well does the image on the right capture the identity of these concepts? (Scale: 1 = Not at all, 5 = Very much). Our method achieves significantly higher ratings for identity alignment compared to other methods, indicating its superior ability to maintain concept identity in composite images, see Tab. 2.
@@
Merging Time. In terms of time efficiency, our method demonstrates significant advantages. It takes only 5 minutes to combine 12 LoRA models. In contrast, [24] requires fine-tuning each LoRA model from scratch, each taking approximately 10-15 minutes. While the actual merging process of [24] takes only 1 second, it requires prior fine-tuning, which adds up to a total of 120 minutes to merge them. Meanwhile, Mix-of-Show requires 15 minutes to merge the models. After merging, generating the images takes approximately 10 seconds for all methods. This comparison clearly shows that our method is substantially faster than other methods in merging models. Once LoRACLR merges LoRA models into a unified model, it can generate composite images without any further need for retraining the individual concepts or access to original training data.
4.3 Ablation Study
We conduct ablation studies to evaluate the impact of key parameters. By exploring the effects of margin, $\lambda_{\text{delta}}$, and concept count, we identify optimal settings for robust identity preservation and coherence in complex compositions.
Impact of Margin and $\lambda_{\text{delta}}$. As shown in Fig. 8, we explore different values of margin and $\lambda_{\text{delta}}$. The results demonstrate that our method achieves robust identity preservation and visual coherence with margin values around 0.25–0.5 and $\lambda_{\text{delta}}$ set at 0.001. Higher values for either parameter lead to diminished performance in maintaining individual identity and prompt coherence.
Effect of # Concepts. Qualitative results in Fig. 8 further demonstrate that our model maintains identity and visual coherence even with complex multi-concept compositions, highlighting its scalability and robustness.
@@ 4.3 Ablation Study
5 Limitation and Societal Impact
Our method uses a novel contrastive learning objective to merge pre-trained LoRA models and achieves successful results in multi-subject image synthesis. However, similar to related work [9, 24], its performance is inherently tied to the capabilities of the underlying LoRA models provided as input. Therefore, the success of LoRACLR in generating coherent and high-quality images depends on the robustness and adaptability of these initial models. This dependency highlights the importance of using well-trained, versatile LoRA models to ensure optimal outcomes. Moreover, given that our method enables sophisticated composition capabilities, it is crucial to consider the potential for misuse, such as in creating deepfakes. Thus, we advocate for the careful use of our method to prevent such applications, promote ethical use, and ensure that advancements in image synthesis contribute positively to technology and society.
6 Conclusion
We introduce LoRACLR, a novel approach that merges multiple LoRA models using a contrastive learning objective. Our method preserves the distinct identities of concepts while enabling the creation of composite images from multiple subjects. LoRACLR is designed to be compatible with any existing LoRA model, ensuring seamless integration with future, more advanced models. Operating as a post-training method, LoRACLR requires only a one-time merging process and can be used with any prompts, utilizing community-available LoRA models. Unlike existing methods that require individual fine-tuning of each LoRA model, our approach offers a more efficient and flexible solution. The effectiveness of LoRACLR is demonstrated through extensive testing, where it shows superior visual quality and coherence compared to other methods.

### v2

@@ 2 Related Work
Merging Multiple Concepts. Combining LoRAs for style and subject control remains an open research challenge. Current methods for multi-concept synthesis often face notable limitations. Weighted summation [34] is a simple approach but suffers from feature interference. AnyDoor [3] enables zero-shot object replacement and composition without per-instance fine-tuning. Mix-of-Show [10] requires specialized Embedding-Decomposed LoRAs (ED-LoRAs) for each concept, which limits compatibility with standard LoRAs. ZipLoRA [37] can merge style and content LoRAs but struggles when multiple content LoRAs are needed. OMG [20] depends on off-the-shelf segmentation to isolate subjects, making its performance highly dependent on segmentation accuracy and the model’s ability to generate multiple objects. Recent advancements have focused on merging multiple specialized models into a unified generative framework. Mix-of-Show [10] effectively merges models using ED-LoRAs but requires access to the original training data, which limits compatibility with community LoRAs, such as those available on platforms like civit.ai [4]. Orthogonal Adaptation [26] introduces constraints to separate attributes across LoRAs, reducing interference; however, it increases training complexity by directly modifying the fine-tuning process, which also requires access to the original data.
@@ 3.2 LoRACLR
LoRACLR’s merging process relies on contrastive loss to ensure compatibility across independently trained LoRA models. Given a pre-trained LoRA $V_{i}$, we first create pairs of input and output features, denoted as $X_{i}$ and $Y_{i}$, respectively, see Fig. 2 (a). Meanwhile, predicted features $\hat{Y_{i}}$ represent the merged model’s output for the same input features, $X_{i}$, i.e., $\hat{Y_{i}}=(W+\Delta W)X_{i}$. The intuition behind our contrastive objective is as follows: positive pairs generated by the same LoRA models should attract, while negative pairs generated by the different LoRAs should repel, see Fig. 2 (b). The contrastive loss objective is defined as:
@@
where $d_{n,i}$ is the negative distance, computed as the minimum distance between the output features $Y_{i}$ for a given concept and the predicted features $\hat{Y}_{j}$ of unrelated concepts, with $j\neq i$. This contrastive objective keeps unrelated concepts distinct while aligning each concept’s output features with the merged model’s predictions, with negative pair loss $d_{n}$ further preventing concept interference and ensuring distinct representations for similar concepts.
Delta-Based Merging.
LoRACLR uses an additive delta, $\Delta W$, to merge LoRA models without altering the base weights directly. Initialized to zero, we learn $\Delta W$ to adjust the pre-trained weights, preserving the integrity of each model while ensuring compatibility. An L2 regularization term is applied to $\Delta W$ to limit its magnitude, ensuring sparsity and minimal adjustments. The optimization objective combines the contrastive merging loss, $\mathcal{L}_{\text{contrastive}}$ and $\mathcal{L}_{\text{delta}}$:
@@ 4 Experiments
Implementation Details.
In all our experiments, we use the Stable Diffusion model [31] with the ChilloutMix checkpoint [^1] https://huggingface.co/windwhinny/chilloutmix for its high-quality image generation capability. Our approach leverages pre-trained LoRA models, eliminating the need to train individual LoRA models from scratch and significantly reducing computational overhead. During the merging process, we use a learning rate of $1e^{-4}$, a margin parameter $m$ of 0.5 to control separation between positive and negative pairs during contrastive learning, and a regularization coefficient $\lambda_{\text{delta}}$ of 0.001 to ensure sparsity in the learned delta. All experiments, including model outputs, and processing, were conducted on an NVIDIA A100 GPU at Virginia Tech. Combining 12 concepts is highly efficient, typically taking about 5 minutes, which makes our approach scalable and practical for real-world applications that require seamless integration of multiple concepts. For our experiments, we utilize pre-trained single-concept LoRA and/or ED-LoRA models as the starting point.
Baselines.
We evaluate our approach against several state-of-the-art baselines in multi subject customization, including: DB-LoRA [32], $\mathcal{P}+$ [41], Custom Diffusion [21], Mix-of-Show [10], and Orthogonal Adaptation (OA) [26]. For DB-LoRA, Federated Averaging (FedAvg) is used to merge models. Custom Diffusion utilizes its optimization-based method for merging, while Mix-of-Show merges using gradient fusion. Orthogonal Adaptation introduces orthogonal transformations to mitigate concept interference during merging. In the case of $\mathcal{P}+$, no fine-tuning of model weights is performed; instead, merging is achieved by directly querying each concept’s token embedding.
Datasets & Metrics.
For all evaluations, we use the same experimental setup and dataset proposed by [26], which includes 12 concept identities, each represented by 16 distinct images of the target concept across various contexts. Following prior work [26, 10], we evaluate our method using three key metrics: Text Alignment, Image Alignment, and Identity Alignment. Text Alignment measures text-image similarity using the CLIP [28] model to ensure that generated images match the input prompts. Image Alignment evaluates the similarity between generated and reference images in the CLIP feature space. Identity Alignment uses the ArcFace [5] model to assess how accurately the target human identity is preserved in the generated images.
4.1 Qualitative Results
We demonstrate the qualitative results of our method for both single and multiple subjects. Given a set of subjects, each represented by a personalized model—for instance, celebrities like Margot Robbie and Taylor Swift (see Fig. 3)—and a text prompt such as ‘<Margot> and <Taylor> in an ancient library with towering shelves,’ our objective is to generate a composite image that integrates these subjects according to the provided text prompt. Using 12 subjects identified by [26], we first merge them into a unified model through our novel contrastive-based objective, applied consistently across all experiments. Unlike methods that require individual fine-tuning for each concept [26], our approach can work with pre-trained models such as LoRA or ED-LoRA models.
Single Concepts.
We first demonstrate our method’s ability to preserve individual identities. Figure 3 (bottom row) confirms that our approach maintains the integrity of each identity for single concepts. This capability extends to diverse settings. For example, the “<CONCEPT> Cyberpunk style’ shows subjects portrayed as game-like characters, as seen in the fourth images of the bottom row in Fig. 3.
Figure 5: Non-human subject & Complex Composition. (a) Our method effectively integrates diverse concepts—including animals, objects (e.g., tables, chairs, vases), and landmarks (e.g., pyramids, rocks)—into visually coherent and semantically meaningful scenes. (b) LoRACLR also enables complex interactions between concepts, such as multiple subjects engaging in contextual activities.
Multiple Concepts.
Figure 3 showcases images generated with our method using varying numbers of concepts—6, 5, 4, 3, and 2. These visuals demonstrate that our method not only accurately captures each individual identity but also generates composite images guided by the text prompts. Notably, our approach excels by utilizing a single merged model to generate a wide range of concepts across diverse prompts, eliminating the need to train separate models for different concept counts. We also note that the original LoRA models for individual subjects are trained on various images of celebrities who may have different hairstyles or colors. As a result, variations such as different hairstyles or colors can appear even within generations of the same celebrity model (e.g., different generations of a Taylor Swift LoRA might show varying hair colors or styles, reflecting her real-life changes). These variations are not inconsistencies in our model but rather reflect the diversity inherent in the original LoRA models. Importantly, despite these variations, our model maintains fidelity and identity to the celebrities’ faces and ensures that details like hairstyles do not mix across different subjects.
Qualitative Comparisons.
In Fig. 4, we compare our method against several state-of-the-art approaches, including OA, MoS, and $\mathcal{P}+$, across various concepts. Figure 4 illustrate that our method successfully preserves individual identities without crossovers, whereas other methods encounter issues. For instance, OA [26] inadvertently transfers features, such as the hair of female characters, to others, while MoS [10] and $\mathcal{P}+$ [41] struggle with precise identity depiction. Additionally, while these models perform adequately with two or three concepts, their ability to accurately represent more concepts diminishes as the concept count increases. For instance, in scenes with six concepts, see Fig. 4 first section, methods other than LoRACLR fail to preserve the identity of Messi, highlighting their limitations as the number of concepts increases.
Style LoRA Integration.
To showcase the flexibility of our approach, we integrate style-specific LoRA models to generate scenes combining conceptual and stylistic elements, enabling outputs in styles like comic art or oil painting. As shown in Fig. 7, prompts such as “…in a city, in comic style” capture the vibrant aesthetic of comic art, while “…in a garden, holding flowers, in oil painting style” reflect the textured quality of oil painting. Complex scenes like “…in a castle, signing papers, in oil painting style” illustrate the model’s ability to maintain content coherence alongside stylistic adaptation. These results demonstrate our method’s efficiency in preserving content accuracy and achieving high stylistic fidelity, making it practical for creative and artistic workflows.
@@ 4.1 Qualitative Results
Figure 6: Quantitative Results on Number of Concepts. Text alignment, image alignment, and identity preservation scores remain high as merged concepts increase. Our method maintains identity and prompt adherence. Dots indicate baseline metrics for each LoRA model pre-merging for performance reference.
@@
Figure 7: Style LoRAs. Our method combines styles like comic art and oil painting into multi-subject scenes, ensuring stylistic fidelity and content coherence and showcasing its flexibility.
Non-Human Examples.
@@
We present examples involving non-human concepts, including animals, objects, and monuments. Our model generalizes seamlessly to these scenarios while maintaining high-quality outputs and stylistic consistency. As shown in Fig. 5, our method effectively integrates diverse concepts into cohesive scenes. It excels with objects like tables, chairs, and vases, accurately capturing their distinct textures. Additionally, it demonstrates robustness with monumental elements such as pyramids and rocks, blending them seamlessly with other concepts. These results highlight the adaptability of our method in generating coherent and visually engaging compositions across a wide range of non-human concepts, including animals, objects, and landmarks. This versatility makes it suitable for both creative and practical applications, such as wildlife illustration, interior design, and architectural visualization.
@@ 4.2 Quantitative Results
Effect of Number of Concepts.
Unlike other methods, which struggle to maintain text alignment, image alignment, and identity preservation as the number of combined concepts increases, LoRACLR preserves these metrics. As shown in Fig. 6, our method remains effective even as the number of concepts grows.
User Study.
To supplement the findings in Tab. 1 and Fig. 6, we conduct a user study involving 50 participants on Prolific.com [27]. The study consisted of 40 questions where participants are shown reference images of individual concepts alongside composite images, with varying numbers of concepts, generated by our method and competing methods, presented in a randomized order. Participants are asked to evaluate each image pair based on: Identity Alignment: Given the reference images on the left, how well does the image on the right capture the identity of these concepts? (1 = Not at all, 5 = Very much). Our method achieves significantly higher ratings for identity alignment compared to other methods, indicating its superior ability to maintain concept identity in composite images, see Tab. 2.
@@
Merging Time.
In terms of time efficiency, our method demonstrates significant advantages. It takes only 5 minutes to combine 12 LoRA models. In contrast, [26] requires fine-tuning each LoRA model from scratch, each taking approximately 10-15 minutes. While the actual merging process of [26] takes only 1 second, it requires prior fine-tuning, which adds up to a total of 120 minutes to merge them. Meanwhile, Mix-of-Show requires 15 minutes to merge the models. After merging, generating the images takes approximately 10 seconds for all methods. This comparison clearly shows that our method is substantially faster than other methods in merging models. Once LoRACLR merges LoRA models into a unified model, it can generate composite images without any further need for retraining the individual concepts or access to original training data.
Ablation Study.
We conduct ablation studies to evaluate the impact of key parameters such as the effects of margin, $\lambda_{\text{delta}}$, and concept count as follows:
(i) Impact of Margin and $\lambda_{\text{delta}}$.
As shown in Fig. 8, we explore different values of margin and $\lambda_{\text{delta}}$. The results demonstrate that our method achieves robust identity preservation and visual coherence with margin values around 0.25–0.5 and $\lambda_{\text{delta}}$ set at 0.001. Higher values for either parameter lead to diminished performance in maintaining individual identity and prompt coherence.
(ii) Effect of Number of Concepts.
Qualitative results in Fig. 8 further demonstrate that our model maintains identity and visual coherence even with complex multi-concept compositions, highlighting its scalability and robustness.
@@
5 Discussion
Limitations.
Similar to [10, 26], the performance of our method is inherently tied to the capabilities of the underlying LoRA models. Therefore, the success of LoRACLR in generating coherent and high-quality images depends on the robustness of these initial models. Moreover, given that our method enables sophisticated composition capabilities, it is crucial to consider the potential for misuse, such as deepfakes. Thus, we advocate for the careful use of our method to prevent such applications and promote ethical use.
Conclusion.
We introduce LoRACLR, a novel approach that merges multiple LoRA models using a contrastive learning objective. Our method preserves the distinct identities of concepts while enabling the creation of composite images from multiple subjects. LoRACLR is designed to be compatible with any existing LoRA model, ensuring seamless integration with future, more advanced models. Operating as a post-training method, LoRACLR requires only a one-time merging process and can be used with any prompts, utilizing community-available LoRA models.